OpenStudy (anonymous):

an airplane travels at 950km/h.How long does it take to travel 1.00km?

OpenStudy (anonymous):

speed=distance/time Here speed=950km/h distance=1 km time=distance/speed

OpenStudy (anonymous):

velocity of airplane = 950 km/h = 263.888 m/s distance= 1000 m time = distance/time= =1000/263.88= =3.789 seconds